If the average velocity of a free falling body is numerically equal to half the acceleration due to gravity. Find the velocity of the body as it reaches the ground

Grade:9

2 Answers

Khimraj
5 years ago
average velocity is given by  (1/T)\int_{0}^{T}vdt = g/2
(1/T)\int_{0}^{T}gtdt= g/2
gT/2 = g/2
So T = 1 sec.
Then v = gt = 10*1 = 10 m/s.
Shamnas
5 years ago
A freely falling body has average velocityV(av.) = (1/2). numerical value of g in m/ssuppose it travelled for t secsso velocity of striking the ground v = u + g.t = g.tif it has travelled a distance h in t secondsh= (1/2) g.t^2but the same distance it must have travelled with average velocity V(av.)h = V(av.). t =(g/2).t = (1/2) g.t^2therefore t= 1 sso, v = g.1 m/sthe velocity with which it strikes the ground = g m/s